OpenStudy (anonymous):

hey guys can you please help with this question... i'll post it up in a sec...

OpenStudy (anonymous):

\[\log_{a} (x+y)= \log_{a}x + \log_{a}y \] Rewrite this relation in index form (that is without using logarithms)...

OpenStudy (anonymous):

What is your question can you explain it more to me??

OpenStudy (anonymous):

is it : \[x + y = xy\]

OpenStudy (anonymous):

no you just have to find the relation between the two formulas by using indecies and no logarithms in the final answer.

OpenStudy (anonymous):

What do you mean by index form??

OpenStudy (richyw):

\[\text{use } a^xa^y=a^{x+y}\]

OpenStudy (anonymous):

@waterineyes yes your answer was correct... can you please show me how you got it...

OpenStudy (anonymous):

See use the formula on right hand side: \[Log_x(a) + Log_x(b) = Log_x(a b)\] So, \[Log_a(x + y) = Log_a(xy)\] Taking Anti Logarithm both the sides, \[x + y = xy\]

OpenStudy (anonymous):

so haha anti log? i get it sorta...

OpenStudy (anonymous):

Anti Log means reverse of log like anti derivative etc etc..

OpenStudy (anonymous):

ok thanks...

OpenStudy (anonymous):

Welcome dear...